Calculate


SS,


variance, and standard deviation for the following population of


N


= 5 scores:


2, 13, 4, 10, 6.




1
Expert's answer
2021-09-12T23:52:07-0400

The mean is


"\mu=\bar{x}=\dfrac{2+13+4+10+6}{5}=7"

Then


"SS(sum\ of\ squares)=(2-7)^2+(13-7)^2"

"+(4-7)^2+(10-7)^2+(6-7)^2=80"

Variance


"Var(X)=\sigma^2=\dfrac{SS}{n}=\dfrac{80}{5}=16"

Standard deviation

"\sigma=\sqrt{\sigma^2}=\sqrt{16}=4"